# post-last exit time process and its application to Loss-Given-Default Distribution

### v3

@@ Abstract.
We study a linear diffusion process after its last exit time from a certain regular point. Rather than treating the process as newly born at the last exit time, we view the whole path and separate the original process before and after the last exit time. This enables us not only to identify the transition semigroup, boundary behavior, entrance law, and reverse of the post-last exit time process, but also to establish a financial model for estimating the loss-given-default distribution of corporate debt (an all-time important open problem).
${}^{1}$hone: +81-75-753-3430. ${}^{2}$hone: +81-75-753-3429.
This version: March 7, 2024. An earlier version of this paper has been circulated under the title “A New Approach to Estimating Loss-Given-Default Distribution”. The first author is in part supported by Grant-in-Aid for Scientific Research (C) No.23K01467, Japan Society for the Promotion of Science. The second author is in part supported by JSPS KAKENHI Grant-in-Aid for Early-Career Scientists No.23K12501.
Keywords: linear diffusion; last exit time; excursion; loss given default distribution
@@ 1. Introduction
Let $X=\{X_{t},t\geq 0\}$ be a linear diffusion. We study the process after its last exit time from a certain regular point: $(X_{t})_{t>L_{\alpha}}$ where $L_{\alpha}$ is the last exit time from state $\{\alpha\}$. We call this $(X_{t})_{t>L_{\alpha}}$ the post-last exit time process. The last exit time is an important subject in the probability literature due to being closely related to the concepts of transience/recurrence, Doob’s h-transform, time-reversed processes, and the Martin boundary theory. For these interrelated topics, we refer the reader to Doob (1957), Nagasawa (1964), Kunita and Watanabe (1966), Salminen (1984), Rogers and Williams (1994a), Borodin and Salminen (2002), Chung and Walsh (2004), Revuz and Yor (2005) as well as the literature referenced therein.
While the last exit time is not a stopping time, it has been utilized widely in financial applications as discussed in Nikeghbali and Platen (2013). These applications include credit risk analysis (Egami and Kevkhishvili (2020)), valuation of defaultable claims (Elliott et al. (2000), Jeanblanc and Rutkowski (2000), Coculescu and Nikeghbali (2012), Jeanblanc et al. (2009)), option valuation (Profeta et al. (2010), Cheridito et al. (2012)), and insider trading (Imkeller (2002)).
@@
The contribution of this paper is twofold: mathematical and practical. We study the post-last exit time process in Section 2. It is, to our knowledge, first studied by Meyer et al. (1972) where the process $(X_{L_{\alpha}+t},\mathcal{F}_{L_{\alpha}+t})$ is proved to be a strong Markov process and its transition semigroup is identified. While it is treated as a newborn process at time $L_{\alpha}$ in Meyer et al. (1972), we treat the post-last exit time process in a continuum of the original process $X$: we look at the entire path. This approach allows us to identify and prove, as to the post-last exit time process, its transition semigroup (Proposition 2.1), the state $\{\alpha\}$ being an entrance boundary (Proposition 2.2), its entrance law (Proposition 2.3), and its reversed process (Proposition 2.4). These results are summarized in Table 1, which is rather a comprehensive study.
@@
As a practical contribution, we have a significant financial application: estimating loss-given-default (LGD, hereafter) distribution of corporate debt. This subject has long been an open issue. The model is set up in Section 3 and an empirical study is done in Section 4. The essential points are as follows. While the firm-value approach for credit risk analysis proposed in Merton (1974) is a fundamental method (as is the intensity-based approach) in estimating the default probability, one has some difficulties in calculating the LGD under this approach. For example, if the firm’s default is defined as the first hitting time of the firm value to a certain point, say $c$, then the value upon default is necessarily equal to the deterministic number $c$. To overcome this discrepancy, we consider the last exit time of the firm value (measured by its leverage ratio) to another threshold level $\alpha>c$. This approach has a useful feature. Since the last exit time is not a stopping time, it is convenient when modeling the reality: during the period while the observer does not know whether the firm value would recover to $\alpha$, a default occurs as a surprise. The post-last exit time process well captures this reality and is suitable for estimating LGD distribution. In fact, we compute LGD distribution through the intensity-based approach applied to the post-last exit time process. In this modeling, due to the mathematical contribution mentioned above, we can deal with the firm value process consistently and continuously: the firm-value approach before time $L_{\alpha}$ and intensity-based approach after $L_{\alpha}$.
@@
With minimal and standard assumptions, the model to be presented allows us to obtain the LGD distribution implied in the current CDS market: we compute (1) the model-driven LGD distribution and (2) CDS spreads implied by the estimated LGD distribution. When checking the consistency of the model-implied CDS spreads with the market quotes, we compare “spreads per 1% loss given default” because the market CDS spreads are quoted under the assumption that the LGD is fixed at 60% for all firms. In the end, we emphasize that one should not confuse our loss right upon the default with the final loss after liquidation and/or bankruptcy procedure. For our analysis, the risk-free rate was obtained from the website of the U.S. Department of the Treasury. The remaining data was obtained from Refinitiv Eikon. We refer the interested reader to an extended literature review of LGD distribution in Appendix A.3.
2. A study of post-last exit time process
Fix a probability space $(\Omega,\mathcal{H},\mathbb{P})$ and a filtration $\mathcal{F}=(\mathcal{F}_{t})_{t\geq 0}$. Let $X=\{\omega(t),t\geq 0;\mathbb{P}_{x}\}$ be a linear regular canonical diffusion starting at $x\in\mathbb{R}$: that is,
$\omega(t)=X_{t}(\omega),\quad t\geq 0.$
@@ 2. A study of post-last exit time process
Its state space is given by $\mathcal{I}=(\ell,r)\subset\mathbb{R}$. Let the last exit time from a regular point $\alpha\in\mathcal{I}$ be denoted by
$L_{\alpha}:=\sup\{t:\omega(t)=\alpha\}$
@@
with $\sup\emptyset=0$. The last exit time is an example of co-optional time satisfying
$L_{\alpha}\circ\theta_{t}=(L_{\alpha}-t)^{+},\quad t\geq 0,$ (2.1)
@@
$\forall t:(X_{s}\circ\theta_{t})(\omega)=X_{s}(\theta_{t}\omega)=X_{t+s}(\omega).$ (2.2)
2.1. Elements of Linear Diffusion
We refer the reader to Chapter II in Borodin and Salminen (2002) for the basic facts regarding linear diffusions. The scale function and the speed measure of $X$ are denoted by $s(\cdot)$ and $m(\cdot)$, respectively. We assume that $s$ and $m$ are absolutely continuous with respect to the Lebesgue measure and have smooth derivatives:
@@ 2.1. Elements of Linear Diffusion
where $m$ and $s^{\prime}$ are continuous and positive. If $X$ hits $\ell$ or $r$, it is killed and transferred immediately to the cemetery $\Delta\notin\mathcal{I}$. The lifetime of $X$ is given by
$\zeta=\inf\{t:\omega(t-)=\ell\;\text{or}\;r\}.$
We assume that $X$ is transient. In this paper, we call a diffusion transient (following Salminen (1984)) if
$\forall x\in\mathcal{I},\;\text{$A$ is compact in $\mathcal{B}(\mathcal{I})$: }\mathbb{P}_{x}(L_{\alpha}<\infty)=1.$
The transience in the above sense is equivalent to one or both of the boundaries being attracting; that is, $s(\ell)>-\infty$ and/or $s(r)<+\infty$. It is possible to obtain transient diffusion from originally recurrent diffusion by including an absorbing boundary in its state space. Such approach is often employed in engineering, finance, and other scientific fields when handling real-life problems. Thus, transient diffusions have a wide range of applications. To obtain concrete results, we set a specific assumption:
@@
$\mathbb{P}\left(X_{t+s}\in A\mid\mathcal{F}_{s}\right)=P_{t}(X_{s},A),\quad\forall A\in\mathcal{B}(\mathcal{I})\quad a.s.$
@@
The transition density $p$ may be taken to be jointly continuous in all variables and symmetric such that $p(t;x,y)=p(t;y,x)$.
The Laplace transform of the hitting time $H_{z}:=\inf\{t:\omega(t)=z\}$ for $z\in\mathcal{I}$ is given by
$\mathbb{E}_{x}\left[e^{-qH_{z}}\right]=\begin{cases}\frac{\psi_{q}(x)}{\psi_{q}(z)},\quad x\leq z,\\ \frac{\phi_{q}(x)}{\phi_{q}(z)},\quad x\geq z,\end{cases}$ (2.3)
where the continuous positive functions $\psi_{q}$ and $\phi_{q}$ represent linearly independent solutions of the ODE $\mathcal{G}f=qf$ with $q>0$. Here the function $\psi_{q}$ is increasing while $\phi_{q}$ is decreasing. Once the boundary conditions at $\ell$ and $r$ are specified, these functions are unique up to a multiplicative constant. The Green function is defined as
@@
$G_{q}(x,y):=\begin{cases}\frac{\psi_{q}(x)\phi_{q}(y)}{w_{q}},\quad x\leq y,\\ \frac{\psi_{q}(y)\phi_{q}(x)}{w_{q}},\quad x\geq y\end{cases}$ (2.4)

### v4

# Loss-Given-Default Modeling by post-last passage time process

@@ Abstract.
This study proposes a stochastic model for loss-given-default (LGD) which provides the LGD distribution based on credit market and company-specific financial conditions. The model utilizes last passage time of a linear diffusion (representing firm value) to a certain threshold point, after which default occurs as a surprising event. By treating the post-last passage time process in a continuum of the original process, we are able to use firm-value approach before and intensity-based approach after the last passage time, leading to a hybrid model. Under minimal and standard assumptions, we obtain the distributions of default time and LGD explicitly. We provide a computationally simple estimation procedure and real-world examples of estimated LGD distribution implied in CDS market.
1Phone: +81-75-753-3430. 2Phone: +81-11-706-3174.
This version: November 2, 2025. An earlier version of this paper has been circulated under the title “A New Approach to Estimating Loss-Given-Default Distribution” or “Post-last exit time process and its application to loss-given-default distribution”. The first author is in part supported by Grant-in-Aid for Scientific Research (C) No.23K01467, Japan Society for the Promotion of Science. The second author is in part supported by JSPS KAKENHI Grant-in-Aid for Early-Career Scientists No.23K12501.
Keywords: loss given default distribution; credit risk; asset-to-debt ratio; linear diffusion; last passage time
@@ 1. Introduction
This study is novel in that we have no assumptions on types of distribution or dynamics of loss-given-default (hereafter, LGD); yet, we are able to estimate the LGD distribution density, not merely an expected value. Our LGD distribution is implied in the CDS (credit default swap) market. As Gambetti et al. (2019) pointed out, general economic uncertainty has a considerable impact on company-specific LGD distribution. Considering this economic uncertainty is reflected in the price formation of the CDS market, a model that incorporates the market information should be of use and importance. Furthermore, an estimate of whole distribution density is appreciated from banks’ and investors’ credit risk management point of view.
There exists a large body of literature regarding the LGD distribution because the information about this distribution is important for credit risk management. One example of its use is the estimation of loan-loss reserves in the banking industry. The LGD distribution should reflect credit market conditions as well as company-specific financial conditions. However, as Doshi et al. (2018) and many other studies point out, it is a standard practice to assume, irrespective of company names, a constant recovery rate (approximately 40%) of the debt upon default. This $60\%$ loss rate is arbitrarily determined, possibly from the empirical distribution. Since the market data is usually quoted based on the predetermined constant loss rate (60%) for all firms, it is hard to estimate the LGD distribution by just observing the market information such as CDS spreads, default time distribution, and default probability. In contrast to the standard assumption of the constant loss rate, the empirical literature has documented the evidence of time-varying realized loss rates. Hence we cannot overemphasize the effectiveness of a model that provides the LGD distribution implied in the credit market and company-specific financial conditions.
This subject has long been an open issue. As the empirical literature has documented the evidence of time-varying realized loss rates, it is important to build a stochastic model for LGD. Gambetti et al. (2019) analyzes determinants of recovery rate distributions and finds economic uncertainty to be the most important systematic determinant of the mean and dispersion of the recovery rate distribution. For their analysis, the authors use post-default bond prices of 1831 American corporate defaults during 1990-2013. Even though recovery rates vary with a firm’s idiosyncratic factors, Gambetti et al. (2019) states that the impact of systematic factors related to economic cycles should not be underestimated. For the summary of studies related to cross-sectional and time variation of recovery rates, we refer the reader to Gambetti et al. (2019).
Finally, Cohen and Costanzino (2017) is an example of a study that extends a structural credit risk model and incorporates stochastic dynamics of the recovery rate. In contrast to our approach which does not introduce an additional recovery risk factor and is based solely on the dynamics of the asset-to-debt ratio process, Cohen and Costanzino (2017) models asset and recovery processes separately as correlated geometric Brownian motions. In their model, the asset risk driver serves as a default trigger and the recovery risk driver determines the amount recovered upon default. The authors explicitly compute the prices of bonds and CDS under this framework. See also Kijima et al. (2009).
@@
We take a completely different approach to modeling LGD distribution: using the post-last passage time process. Let $X=\{X_{t},t\geq 0\}$ be a linear diffusion. We study the process after its last passage time from a certain regular point: $(X_{t})_{t>\lambda_{\alpha}}$ where $\lambda_{\alpha}$ is the last passage time from state $\alpha$. We call this $(X_{t})_{t>\lambda_{\alpha}}$ the post-last passage time process. The post-last passage time process in Section 3.2 is, to our knowledge, first studied by Meyer et al. (1972) where the process $(X_{\lambda_{\alpha}+t},\mathcal{F}_{\lambda_{\alpha}+t})$ is proved to be a strong Markov process and its transition semigroup is identified. While it is treated as a newborn process at time $\lambda_{\alpha}$ in Meyer et al. (1972), we treat the post-last passage time process in a continuum of the original process $X$: we look at the entire path, a contributing factor for the modeling of LGD distribution.
Remark 1.1 (Economic interpretation of last passage time).
@@
The model is discussed in Section 2 and an empirical study is performed in Section 5. The essential rationale of using last passage time is as follows. While the firm-value approach for credit risk analysis proposed in Merton (1974) is a fundamental method (as is the intensity-based approach) in estimating the default probability, one has some difficulties in calculating the LGD under this approach. For example, if the firm’s default is defined as the first hitting time of the firm value to a certain point, say $c$, then the value upon default is necessarily equal to the deterministic number $c$. To overcome this discrepancy, we consider the last passage time of the firm value (measured by its asset-to-debt ratio) to another threshold level $\alpha>c$. First, the fact that the firm value will not recover to level $\alpha$ after time $\lambda_{\alpha}$ is significant since it implies that the company is not creditworthy after $\lambda_{\alpha}$. We shall make use of this observation. Moreover, this approach has a useful feature. Since the last passage time is not a stopping time, it is convenient when modeling the reality: during the period while the observer does not know whether the firm value would recover to $\alpha$, a default occurs as a surprise. Moreover, the post-last passage time process well captures this reality and is suitable for estimating LGD distribution. In fact, we compute LGD distribution through the intensity-based approach applied to the post-last passage time process. In this modeling, due to the mathematical contribution (in Section 3.2) which allows us to look at the entire path of $X$ as mentioned above, we can deal with the firm value process consistently and continuously: the firm-value approach before time $\lambda_{\alpha}$ and intensity-based approach after $\lambda_{\alpha}$. Another advantage of the modeling is that we do not need to detect or predict when the last passage time occurs in the estimation of LGD distribution.
@@
In Section 4, with minimal and standard assumptions, the model to be presented allows us to obtain the LGD distribution implied in the current CDS market: we compute (1) the model-driven LGD distribution and (2) CDS spreads implied by the estimated LGD distribution. The results are the explicit formula for the distribution of default time in Proposition 4.1 and that of LGD in Corollary 4.2. We also include a procedure to make sure whether the estimated LGD reflects the market data. When checking the consistency of the model-implied CDS spreads with the market quotes, we compare “spreads per 1% loss given default” because the market CDS spreads are quoted under the assumption that the LGD is fixed at 60% for all firms. In the end, we emphasize that one should not confuse our loss right upon the default with the final loss after liquidation and/or bankruptcy procedure. For our analysis, the risk-free rates were obtained from the website of the Federal Reserve Bank of St.Louis. The remaining data was obtained from LSEG Workspace.
Section 5 is devoted to empirical study, where we compute three probability densities; in addition to the LGD distribution (Figure 1) of eight companies, we show the distribution of the last passage time $\lambda_{\alpha}$ (Figure 2) and the occupancy time above level $\alpha$ before the process is killed at some levels of $c<\alpha$ (Figure 3). The former is based on the study in Salminen (1984) and Egami and Kevkhishvili (2020), and the latter is a new result derived in Proposition 3.1. All three distributions are, we believe, important and useful for risk management.
The last passage time is an important subject in the probability literature due to being closely related to the concepts of transience/recurrence, Doob’s $h$-transform, time-reversed processes, and the Martin boundary theory. For these interrelated topics, we refer the reader to Doob (1957), Nagasawa (1964), Kunita and Watanabe (1966), Salminen (1984), Rogers and Williams (1994a), Borodin and Salminen (2002), Chung and Walsh (2004), Revuz and Yor (2005) as well as the literature referenced therein. While the last passage time is not a stopping time, it has been utilized widely in financial applications as discussed in Nikeghbali and Platen (2013). These applications include credit risk analysis (Egami and Kevkhishvili, 2020), valuation of defaultable claims (Elliott et al. (2000), Jeanblanc and Rutkowski (2000), Coculescu and Nikeghbali (2012), Jeanblanc et al. (2009)), option valuation (Profeta et al. (2010), Cheridito et al. (2012)), and insider trading (Imkeller, 2002).
2. Loss-given-default Modeling
2.1. Basic idea
The goal is to obtain the LGD distribution implied in the current CDS market and company-specific financial conditions. More specifically, we derive the distribution of the ratio of the firm’s total assets over its total debt upon default. We shall hereafter refer to the dynamics of this ratio as the asset-to-debt ratio process, denoted by $Y$. In our framework, we are focusing on the loss rate of the debt of the whole company, not each individual debt obligation. Based on this corporate-level loss distribution, credit managers can calculate loss rates for individual debt obligations having distinct characteristics; for example, senior or subordinated, and with or without collateral. This is convenient for empirical study because the spreads in the CDS market are quoted in terms of the risk of the corporate name, not that of its specific debt obligation.
@@ 2.1. Basic idea
Fix a probability space $(\Omega,\mathcal{H},\mathbb{P})$ and a filtration $\mathcal{F}=(\mathcal{F}_{t})_{t\geq 0}$. Let $X=\{\omega(t),t\geq 0;\mathbb{P}_{x}\}$ be a linear regular canonical diffusion starting at $x\in\mathbb{R}$: that is, $\omega(t)=X_{t}(\omega),t\geq 0$. Its state space is given by $\mathcal{I}=(\ell,r)\subset\mathbb{R}$. Let the last passage time from a regular point $\alpha\in\mathcal{I}$ be denoted by
$\lambda_{\alpha}:=\sup\{t:\omega(t)=\alpha\}$
@@
with $\sup\emptyset=0$. The last passage time is an example of co-optional time satisfying
$\lambda_{\alpha}\circ\theta_{t}=(\lambda_{\alpha}-t)^{+},\quad t\geq 0,$
@@
$\forall t:(X_{s}\circ\theta_{t})(\omega)=X_{s}(\theta_{t}\omega)=X_{t+s}(\omega).$
In the firm-value (or Merton (1974)) model of credit risk, insolvency occurs upon the asset-to-debt ratio being equal to some level $c$. In this case, if we set $\alpha>c$, then there must exist the last passage time $\lambda_{\alpha}$ recorded by the asset-to-debt ratio process (to be defined in the next Section 2.3). The idea is summarized as follows:
  1. (1)
  2. (2)
    Since $\lambda_{\alpha}$ is not a stopping time, it is appropriate to model the interval, denoted by $\tau$, between $\lambda_{\alpha}$ and default time $\xi$ by the intensity-based model:
    $\xi(\omega):=\lambda_{\alpha}(\omega)+\tau\circ\theta(\lambda_{\alpha}(\omega)).$
    In fact, this well explains the reality that a default occurs surprisingly during the period when non-insider investors do not have enough information about credit quality. The $\lambda_{\alpha}$, not a stopping time, corresponds to the time when creditworthiness becomes hard to grasp, and $\xi$ occurs as a surprise after non-predicting $\tau$ has elapsed.
  3. (3)
    The company’s asset-to-debt ratio at $\xi$ is then computed as the distribution of the value of the asset-to-debt ratio process at time $\xi$, from which the LGD distribution at the corporate level is obtained. For this purpose, we use results from Section 3.2.
We explicitly derive the distribution of default time $\xi$ (Proposition 4.1) and LGD (Corollary 4.2). These are computationally simple formulas. Furthermore, the model check proposed in Section 4.3 inspects whether the model-implied default time and LGD distribution are consistent with the credit market and company-specific financial conditions. Specifically, we check that the CDS spread implied from the estimated LGD distribution is consistent with the quoted spread in the market.
Let us emphasize that we treat the entire process in a continuum starting at time zero, rather than treating the process after its passage time as a newborn. This is important since for practical purposes we wish to observe the process through its lifetime (from time zero to the killing). It is particularly true when we look at a company’s creditworthiness and estimate its LGD distribution.
2.2. Mathematical Setting
Let us return to the general mathematical setup at the beginning of this section for a general diffusion $X$. We refer the reader to Chapter II in Borodin and Salminen (2002) for the basic facts regarding linear diffusions. The scale function and the speed measure of $X$ are denoted by $s(\cdot)$ and $m(\cdot)$, respectively. We assume that $s$ and $m$ are absolutely continuous with respect to the Lebesgue measure and have smooth derivatives:
@@ 2.2. Mathematical Setting
where $m$ and $s^{\prime}$ are continuous and positive. If $X$ hits $\ell$ or $r$, it is killed and transferred immediately to the cemetery $\Delta\notin\mathcal{I}$. We assume that $X$ is transient. In this paper, we call a diffusion transient (following Salminen (1984)) if
$\forall x\in\mathcal{I},\;\text{$A$ is compact in $\mathcal{B}(\mathcal{I})$: }\mathbb{P}_{x}(\lambda_{A}<\infty)=1\;\textnormal{with}\;\lambda_{A}=\sup\{t:\omega(t)\in A\}.$
The transience in the above sense is equivalent to one or both of the boundaries being attracting; that is, $s(\ell)>$-$\infty$ and/or $s(r)<+\infty$. To obtain concrete results, we set a specific assumption:
@@
Hence the process is killed at the lower boundary $\ell$, which is consistent with our asset-to-debt ratio process.
@@
$\mathbb{P}\left(X_{t+s}\in A\mid\mathcal{F}_{s}\right)=P_{t}(X_{s},A),\quad\forall A\in\mathcal{B}(\mathcal{I})\quad\mathbb{P}-a.s.$
@@
The transition density $p$ may be taken to be jointly continuous in all variables and symmetric such that $p(t;x,y)=p(t;y,x)$. Note that the Green function $G(y,z)$ for $y,z\in\mathcal{I}$ has a representation in terms of the scale function:
@@
$G(y,z):=\int_{0}^{\infty}p(t;y,z){\rm d}t=\begin{cases}\lim\limits_{a\downarrow\ell,b\uparrow r}\frac{(s(y)-s(a))(s(b)-s(z))}{s(b)-s(a)},\quad\ell<y\leq z<r,\\ \lim\limits_{a\downarrow\ell,b\uparrow r}\frac{(s(z)-s(a))(s(b)-s(y))}{s(b)-s(a)},\quad\ell<z\leq y<r.\end{cases}$ (2.1)
The Laplace transform of the hitting time $H_{z}:=\inf\{t\geq 0:X_{t}=z\}$ for $z\in\mathcal{I}$ is given by
$\mathbb{E}_{x}\left[e^{-qH_{z}}\right]=\begin{cases}\frac{\phi_{q}(x)}{\phi_{q}(z)},\quad x\geq z,\\ \frac{\psi_{q}(x)}{\psi_{q}(z)},\quad x\leq z,\end{cases}$ (2.2)
where the continuous positive functions $\psi_{q}$ and $\phi_{q}$ represent linearly independent solutions of the ODE $\mathcal{G}f=qf$ with $q>0$. Note that $\psi_{q}$ is increasing while $\phi_{q}$ is decreasing. These functions are unique up to a multiplicative constant, once the boundary conditions at $\ell$ and $r$ are specified. Finally, the q-Green function is defined as
$G_{q}(x,y):=\begin{cases}\frac{\psi_{q}(y)\phi_{q}(x)}{w_{q}},\quad x\geq y,\\ \frac{\psi_{q}(x)\phi_{q}(y)}{w_{q}},\quad x\leq y,\end{cases}$ (2.3)
with the Wronskian $w_{q}:=\psi_{q}^{+}(x)\phi_{q}(x)-\psi_{q}(x)\phi_{q}^{+}(x)=\psi_{q}^{-}(x)\phi_{q}(x)-\psi_{q}(x)\phi_{q}^{-}(x)$. It holds that $G_{q}(x,y)=\int_{0}^{\infty}e^{-qt}p(t;x,y){\rm d}t$ for $x,y\in\mathcal{I}$.
2.3. Asset-to-debt ratio process
Before providing mathematical results necessary for our analysis in Section 3.2 below, we define the asset-to-debt ratio process which plays the central role in our model. Let us assume that the firm value (market value of total assets) $V$ follows a geometric Brownian motion:
${\rm d}V_{t}=\mu V_{t}{\rm d}t+\sigma V_{t}{\rm d}W_{t}$ (2.4)
with $\mu\in\mathbb{R}$ and $\sigma>0$. The debt process is given by
$B_{t}=B_{0}e^{rt},\quad B_{0}\in\mathcal{F}_{0},\quad t\geq 0,$
where $r>0$ denotes a constant rate. As announced above, we define the asset-to-debt ratio process $Y=V/B$ on $(0,+\infty)$:
The parameters $\mu$, $\sigma$ are such that the left boundary 0 is attracting and the right boundary $\infty$ is non-attracting, which ensures that the default occurs eventually with probability one. This complies with Assumption 1. In particular, we assume that $\mu-\frac{1}{2}\sigma^{2}-r<0$. By fixing a certain level $\alpha\in\mathbb{R}_{+}$, we define
$\lambda_{\alpha}:=\sup\{t\geq 0:Y_{t}=\alpha\}.$